# Cubic BeB2: A metastable -type conductive material from first principles

### v1

@@ IV Defect formation and electronic transport
In order to assess the applicability of -BeB2 as a -type semiconductor, we calculate the defect formation energy of several acceptor point defects, as shown in Fig. 4. The intrinsic defects considered are Be vacancy (VBe), and Be antisites (BeB). Extrinsic point defects considered are H or Li substitutions on the Be site (HBe and LiBe). Notably, for all acceptor-like point defects considered apart from HBe, negative defect formation energies are found. While LiBe and BeB both show an ionization energy of approximately 0.1 eV above the VBM, VBe shows a formation energy of eV and stabilizes in the charge state at the VBM. Therefore, the Be vacancy is expected to be the dominant point defect in -BeB2, acting as a shallow acceptor. These negative formation energies are also found in other compounds, typically suggesting very high doping limits [52] and intrinsic conductivity [53]. In this case, we expect -BeB2 to behave intrinsically as a degenerate -type semiconductor, motivating further investigation of hole transport and potential superconductivity.
@@
Due to the anticipated intrinsic -type conductivity of -BeB2, we next evaluate the hole transport properties in the material with the consideration of both hole-phonon scattering and ionized-impurity scattering. We show the dependence of the intrinsic hole mobility on temperature and the dependence of hole mobility on dopant concentration at room temperature (300 K) in Fig. 5. Compared to the intrinsic hole mobility of diamond, the intrinsic hole mobility in -BeB2 is lower by approximately 40%. However, the hole mobility still exceeds many known electronic materials such as Si and GaAs. As -type conductivity is expected, we examine the hole mobilities from the intrinsic regime up to 1022 cm-3. Our results on the carrier-concentration dependence of the hole mobility show that at low doping levels, hole mobility in the material is mainly limited by hole-phonon scattering. However, above cm-3, ionized-impurity scattering start to become the dominate factor, reducing the mobility by 60% at cm-3 and a factor of 5 when reaching high doping regime such as and cm-3. Despite the strong impact of ionized impurity scattering, we find the hole mobilities to remain on the order of 10-100 cm2V-1s-1 even at very high doping levels, i.e., 86 and 19 cm2V-1s-1 at cm-3 and cm-3, respectively. The resulting electrical conductivities are m-1 and m-1 for the two carrier concentrations. Therefore, a high electrical conductivity is expected in -BeB2 due to the low acceptor formation and ionization energies and the high hole mobility.
@@ V Superconductivity
To further investigate the effect of doping on the phonon dispersion, we also adopt a charge-compensated jellium model (JM). Figure S4 [45] compares the phonon dispersion of the undoped system with that of a doped system within the JM for a doping level of cm-3 corresponding to a Fermi level shift of 1.5 eV below the VBM. The JM leads to noticeable changes in the optical phonon branches, particularly at higher frequencies. Notably, the LO-TO splitting vanishes, which can be attributed to the metallic screening arising from partially occupied electronic states. This screening suppresses the long-range dipole-dipole interactions responsible for LO-TO splitting, leading to a softening of the optical phonons, especially near the point. Such phonon softening significantly affects the electron-phonon coupling.
@@ VI Conclusion
We evaluate the electronic and optical properties of -BeB2. The material is found to be an indirect semiconductor with a fundamental gap on the order of 1.5 eV with very low effective masses for hole transport. Defect calculations reveal that Be vacancy, which is a degenerate acceptor-like defect, dominates defect formation in the material, therefore making it intrinsically -type. We further predict a high intrinsic hole mobility of 1,259 cm2s-1V-1 at room temperature, considering both electron-phonon and electron-ionized-impurity scattering. The mobility of -type -BeB2 is found to be mainly limited by electron-ionized-impurity scattering and remains high even for very high doping levels. Furthermore, we find that heavily doped -type -BeB2 is a low-temperature superconductor, exhibiting a of up to 3.55 K based on Eliashberg theory, comparable to other heavily doped covalent semiconductors such as diamond, Si, and SiC.

### v2

@@ IV Defect formation and electronic transport
In order to assess the applicability of -BeB2 as a -type semiconductor, we calculate the defect formation energy for several acceptor point defects as a function of Fermi level and growth conditions, as shown in Fig. 5. Due to the low formation energy of -BeB2, the allowed chemical potential range is narrow ( eV, eV). As a result, including more stable Be-B compounds will not leave a stable range for the chemical potentials of B and Be for -BeB2. In our analysis, we do not consider the formation of other stable Be-B compounds, assuming that the formation of the stable phase is prevented by the growth condition such as epitaxial strain. We focus on the demonstration of representative shallow acceptors, intrinsic and extrinsic ones, to show the potential of the material for -type conductivity. The intrinsic defects considered are Be vacancy (VBe), and Be antisites (BeB). Extrinsic point defects considered are H or Li substitutions on the Be site (HBe and LiBe). Notably, for all acceptor-like point defects considered apart from HBe, negative defect formation energies are found.
@@
While LiBe and BeB both show an ionization energy of approximately 0.1 eV above the VBM, VBe shows a formation energy of eV and stabilizes in the charge state at the VBM. Therefore, the VBe is expected to be the dominant point defect in -BeB2, acting as a shallow acceptor. We note that negative defect formation energies are similarly found in other nitride and oxide compounds, typically suggesting very high doping limits [55] and intrinsic conductivity [56]. Therefore, we expect -BeB2 to be a potential degenerate -type semiconductor, motivating further investigation of hole transport and potential superconductivity.
Interestingly, the thermodynamic stability of the BeB2.75 composition [1, 6], i.e., a Be-deficient polytype compared to -BeB2, suggests a possible relationship between the negative formation energy of Be vacancies with the increased thermodynamic stability of Be-deficient phases. Our results provide a qualitative understanding of this connection by attributing it to the strengthening of nearby B-B bonds upon removal of the Be atom (See the analysis and Fig. S5 in the supplemental information[32]). However, the stoichiometry of BeB2.75 corresponds to an exceptionally high concentration of Be vacancies ( 27%) in the BeB2 stoichiometry. Moreover, although B adopts a tetrahedral geometry in both polytypes, the two phases have significantly different crystal structures (diamond network in -BeB2 versus icosahedral clusters in BeB2.75). Therefore, whether the formation energy of isolated Be vacancies could serve as an indicator for the stability of Be-deficient phases is not immediately obvious and would require further investigations, which are beyond the scope of the present study.
@@
Due to the anticipated intrinsic -type conductivity of -BeB2, we next evaluate the hole transport properties in the material with consideration of both hole-phonon scattering and ionized-impurity scattering. We show the dependence of the intrinsic hole mobility on temperature and the dependence of hole mobility on dopant concentration at room temperature (300 K) in Fig. 6. Compared to the intrinsic hole mobility of diamond, the intrinsic hole mobility in -BeB2 is lower by approximately 40%. However, the hole mobility still exceeds many known electronic materials such as Si and GaAs. As -type conductivity is expected, we examine the hole mobilities from the intrinsic regime up to 1022 cm-3. Our results on the carrier-concentration dependence of the hole mobility show that at low doping levels, hole mobility in the material is mainly limited by hole-phonon scattering. However, above cm-3, ionized-impurity scattering begins to become the dominant factor, reducing mobility by 60% at cm-3 and a factor of 5 when reaching a high doping regime such as and cm-3. Despite the strong impact of ionized-impurity scattering, we find the hole mobilities to remain on the order of 10-100 cm2V-1s-1 even at very high doping levels, i.e., 86 and 19 cm2V-1s-1 at cm-3 and cm-3, respectively. The resulting electrical conductivities are m-1 and m-1 for the two carrier concentrations. Therefore, a high electrical conductivity is expected in -BeB2 due to the low acceptor formation and ionization energies and the high hole mobility.
@@ V Superconductivity
To further investigate the effect of doping on the phonon dispersion, we also adopt a charge-compensated jellium model (JM). Fig. S6 [32] compares the phonon dispersion of the undoped system with that of a doped system within the JM for a doping level of cm-3 corresponding to a Fermi level shift of 1.5 eV below the VBM. The JM leads to noticeable changes in the optical phonon branches, particularly at higher frequencies. Notably, the LO-TO splitting vanishes, which can be attributed to the metallic screening arising from partially occupied electronic states. This screening suppresses the long-range dipole-dipole interactions responsible for LO-TO splitting, leading to a softening of the optical phonons, especially near the point. Such phonon softening significantly affects the electron-phonon coupling.
@@ VI Conclusion
We evaluate the electronic and optical properties of -BeB2. The material is found to be an indirect semiconductor with a fundamental gap on the order of 1.5 eV with very low effective masses for hole transport. Defect calculations reveal that Be vacancy, which is a shallow acceptor defect with negative formation energy, dominates defect formation in the material, therefore making it intrinsically degenerate -type. We further predict a high intrinsic hole mobility of 1,259 cm2s-1V-1 at room temperature, considering both hole-phonon and hole-ionized-impurity scattering. The mobility of -type -BeB2 is found to be mainly limited by hole-ionized-impurity scattering and remains high even for very high doping levels. Furthermore, we find that heavily doped -type -BeB2 is a low-temperature superconductor, exhibiting a of up to 3.55 K based on Eliashberg theory, comparable to other heavily doped covalent semiconductors such as diamond, Si, and SiC.